OpenStudy (anonymous):

N(t) = 10000 (8+t) e^-0.1t I am trying to find the time when the population will be maximized. I know I have to find derivative equal to 0, I am not sure how to go about from there. For my derivative I got: 10000 (-0.1 e^-0.1t) + (10000 (8+t)) (-0.1e^-o.1t))

OpenStudy (kirbykirby):

\[\frac{d}{dt}N(t)=10000\left[\underbrace{(1)}_{f'}\underbrace{e^{-0.1t}}_{g}+\underbrace{(8+t)}_{f}\underbrace{(-0.1)e^{-0.1t}}_{g'}\right]\]

OpenStudy (anonymous):

I don't know how to solve equal to 0 for my derivative. I get a weird answer.

OpenStudy (kirbykirby):

\[10000\left( e^{-0.1t}+(8+t)(-0.1)e^{-0.1t}\right)=0\\ e^{-0.1t}+(8+t)(-0.1)e^{-0.1t}=0\\ e^{-0.1t}\left[ 1+(8+t)(-0.1)\right]=0, \,\,\,\text{by factoring }e^{-0.1t}\\ e^{-0.1t}[1-0.8-0.1t]=0\\\ e^{-0.1t}(0.2-0.1t)=0\] You have a product of 2 factors, either product should result in 0: But be careful, an exponential can never equal 0, so only the \(( 0.2-0.1t)\) factor can equal 0 \(0.2-0.1t=0\) \(0.1t=0.2\) \(t=2\)

OpenStudy (anonymous):

Oh man thank you so much. I was stuck on this question forever. It makes sense!

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

Do you tutor for math?

OpenStudy (kirbykirby):

I have done some tutoring before